OpenStudy (anonymous):

-.5x + 3 = -x+7? A.) 1/2 B.) 8 C.)-8 D.) 2

OpenStudy (anonymous):

-5x+3=-x+7 =>-5x+x=7-3 =>-4x=4 =>x=-1

OpenStudy (anonymous):

its 1/2 not 5. sorry!

OpenStudy (karatechopper):

-.5x + 3 = -x+7 Add -.5x to both sides -.5x + 3 = -x+7 +.5x +.5x -------------- 3= -.5x + 7 Subtract 7 from both sides 3= -.5x + 7 -7 -7 ------------- -4 = -.5x Divide by -.5 on both sides -4/-.5 = -.5x/-.5 -4/-.5=x 8=x

OpenStudy (anonymous):

ohhhhhhh! thank you!!!ok i have a really hard question! would u mind helping?

OpenStudy (karatechopper):

Sure, but I might be on and off a bit ok?

OpenStudy (anonymous):

ok! how do you solve 2(x+6) - 10 = 12+4(x-1) using only words! please help!

OpenStudy (karatechopper):

words?

OpenStudy (karatechopper):

Hm..

OpenStudy (anonymous):

and maybe on the side it in # form so i understand to hahaha

OpenStudy (karatechopper):

Do the numbers have to be written in words to?

OpenStudy (karatechopper):

alright, here is an idea, this time you solve the equation, any regular way you do it. at the end we will do the words

OpenStudy (anonymous):

yes

OpenStudy (karatechopper):

So..this time you try and solve the equation.

OpenStudy (anonymous):

ok!

OpenStudy (karatechopper):

:) Post your work on here then k?

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

2x + 12 - 10 = 12 + 4x - 4 +10 +10 2x + 12 = 22 + 4x - 4 -12 -12 2x = 34 + 4x -4 -4 -4 -2x = 34 -4 -2x = 30 x = -15?

OpenStudy (anonymous):

?

OpenStudy (karatechopper):

Not quite.. Lemme show u my work.\, then try to see where you messed up

OpenStudy (anonymous):

ok

OpenStudy (karatechopper):

2(x+6)-10 = 12+ 4(x-1) Distribute the 2 and 4 on their own sides 2x+12 -10 = 12 +4x -4 Add/Subtract like terms on their own sides of the equal sign 2x + 2 = 8 +4x Subtract 2x from both sides 2x + 2 = 8 +4x -2x 2x -------------------- 2= 8 + 2x Subtract 8 from both sides 2= 8 + 2x -8 -8 -------------- -6= 2x Divide by 2 on both sides -6/2=2x/2 -3 = x

OpenStudy (anonymous):

ohhhh! ok! lol hard part know... gotta put into words!

OpenStudy (karatechopper):

lol yeah! Just type out how you would to the problem. You know how you say it to yourself like this step comes next, add x to y plus z what ever like that!

OpenStudy (anonymous):

ok!